The perimeter a rectangle is 80 inches. If the width is 18 inches . What is the area of the rectangle?
A) 22
B) 324
C)396
D) 6400

Found 2 solutions by rothauserc, nofear:
Answer by rothauserc(4718)

let l be the length of the rectangle
:
2l +2(18) = 80
:
2l = 44
:
l = 22
:
Area of rectangle = 22 * 18 = 396
:
**********************************
Answer is C)396
**********************************
:

Answer by nofear(2)

18x(40-18)=396